### v2

# Polynomial fingerprinting for trees and formulas

Polynomial fingerprinting for trees and formulas
Mihai Prunescu (1) Research Center for Logic, Optimization and Security (LOS),
Faculty of Mathematics and Computer Science,
University of Bucharest, Academiei 14, 010014 Bucharest, Romania(2) Simion Stoilow Institute of Mathematics of the Romanian Academy,
Research unit 5, P. O. Box 1-764, RO-014700 Bucharest, Romania.(3) Institute for Logic and Data Science, Bucharest, Romania mihai.prunescu@imar.ro, mihai.prunescu@gmail.com
Rezumat
To cater to the needs of (Zero Knowledge) proofs for (mathematical) proofs, we describe a method to transform formal sentences in $2\times 2$ matrices over multivariate polynomials with integer coefficients, such that usual proof-steps like modus-ponens or the substitution are easy to compute from the matrices corresponding to the terms or formulas used as arguments. By evaluating the polynomial variables in random elements of a suitably chosen finite field, the proof is replaced by a numeric sequence. Only the values corresponding to the axioms have to be computed from scratch. The values corresponding to derived formulas are computed from the values corresponding to their ancestors by applying the homomorphic properties. On such sequences, various Zero Knowledge methods can be applied.
1 Motivation
In the remarkable White Paper [5] published by the company Pi Squared Inc. , the following idea is outlined. In order to certificate the correctness of some computation, as for example a block-chain transaction, one can proceed as follows:
  • One proves that the given algorithm, with the given input data, must lead to the concrete output obtained by the computation.
  • The resulting mathematical proof, as also other mathematical proofs in general, are lenghy and tedious. So as they are, they cannot serve as correctness certficate, neither by their length, nor by the time necessary to verify them.
  • At this point one can use Zero Knowledge Proofs of mathematical proofs, in order to reduce their correctness to checking a much smaller certificate. The correctness of this certificate implies with great probability $1-\varepsilon$ the correctness of the mathematical proof. Blum [1] was possibly the first author to express the idea to secure a mathematical proof by a verifiable Zero Knowledge procedure.
  • One can transform the proof in a sequence of numbers, and the most convenient numbers should be the elements of a large finite field. Then one can apply some modern method of Zero Knowledge Proof, instead of the pioneer methods of Blum [1].
  • The lines above outline the concept Proof of Proof, symbolically $\pi(\pi)$. This also explains the acronymous Pi Squared used by the company Pi Squared Inc.
  • It would be even better if this encoding would be homomorphic for usual proof steps, like modus ponens or substitution of variables with terms. This way, one would apply the difficult (time-consuming) encoding only for initial steps, like invocation of axioms or tautologies, while the encodings of other proof steps would be computed from already existing encodings via homomorphic rules.
  • Moreover, if proof-steps like modus-ponens and substitution, become homomorphic, they could be emulated by arithmetic operations. So, various segments of proof can be modeled by arithmetic circuits. Doing so, one can apply Zero Knowledge Proof methods to certify the correctness of the computations done by some arithmetic circuits, like the CIRCOM method, [2].
2 The principle
The following strategy crystalizes:
  1. 1.
    To each formula $\varphi$ corresponds a field element (or a vector consisting of field elements) $V(\varphi)$. Also, to every well-formed term $t$, corresponds a field element (or a vector consisting of field elements) $V(t)$.
  2. 2.
    The vector $V(\varphi)$, respectively $V(t)$, contains a sub-vector $[[\varphi]]$, respectively $[[t]]$, which directly encodes $\varphi$, respectively $t$. The other elements arising in this vector are useful for performing the substitution. Any of the subvector is an auxiliar value to perform the substitution of the variable $x_{i}$ with some other term. So the fingerprint looks like:
    $V(\varphi)=([[\varphi]],[[\varphi]]_{x_{1}},\dots,[[\varphi]]_{x_{k}}).$
  3. 3.
    There is an arithmetic term $MP(a,b)$ with the following property. If a formula $\varphi_{3}$ is the result of applying the rule modus ponens to formulas $\varphi_{1}$ and $\varphi_{2}$, then $V(\varphi_{3})=MP(V(\varphi_{1}),V(\varphi_{2}))$.
  4. 4.
    There is an arithmetic term $Subst_{y}(a,b)$ with the following property. If a formula $\varphi_{3}$ is the result of the substitution of the variable $y$ occurring in the formula $\varphi_{1}$ with the formula $\varphi_{2}$, then $V(\varphi_{3})=Subst_{y}(V(\varphi_{1}),V(\varphi_{2}))$. Also, if a formula $\varphi_{3}$ is the result of the substitution of the variable $y$ occurring in the formula $\varphi_{1}$ with the term $t$, then $V(\varphi_{3})=Subst_{y}(V(\varphi_{1}),V(t))$.
We call the vactor $V(\varphi)$ the fingerprint of $\varphi$.
The definition above is for the time being just a declaration of intentions. These intentions must be completed with some other conditions, as follows:
  • The algorithm should not depend on the particular finite field $\mathbb{F}_{p}$ which is chosen.
  • Different formulas must map on different elements, at least if the characteristic $p$ is sufficiently large, as often as possible. The fingerprinting must be a hash-function.
  • For axioms $\varphi$, the computation of $V(\varphi)$ is fast.
  • For given values of $V(\varphi_{1})$ and $V(\varphi_{2})$, the computation of
    $MP(V(\varphi_{1}),V(\varphi_{2}))$
    is fast.
  • For given values of $V(\varphi_{1})$ and $V(\varphi_{2})$, the computation of
    $Subst_{y}(V(\varphi_{1}),V(\varphi_{2}))$
    is fast.
A way we can solve this problem: We consider a non-commutative ring of $2\times 2$ matrices over the ring $R=\mathbb{Z}[X_{1},X_{2},\dots]$, which is the ring of polynomials with infinitely many variables. We make a first correspondence between formula and terms:
$\varphi\leadsto[\varphi]\in M_{2\times 2}(R),$
respectively
$t\leadsto[t]\in M_{2\times 2}(R),$
satisfying the Unique Encoding Property, which says that some element $A\in M_{2\times 2}(R)$ corresponds to at most one well-formed string, being a formula or a term.
For this encoding, we define the symbolic fingerprint
$F(\varphi)=([\varphi],[\varphi]_{x_{1}},\dots,[\varphi]_{x_{k}}),$
and analogous for terms. The matrices $[\varphi]_{x_{i}}$ are necessary only for computing substitutions, but they must be updated also during modus ponens steps. Please remark that we use the notation $F(\varphi)$ for a finite sequence of matrices over polynomials, while $V(\varphi)$ is a finite sequence of matrices over a finite field $\mathbb{F}$. We obtain $V(\varphi)$ from $F(\varphi)$ by evaluating the polynomial variables in randomly chosen field elements.
The arithmetic terms $MP(a,b)$, respectively $Subst_{y}(a,b)$, work already in the ring $M_{2\times 2}(R)$.
Now, for a random choice of values $X_{1}=r_{1},X_{2}=r_{2},\dots\in\mathbb{F}_{p}$, we evaluate the entries of the $2\times 2$ matrices, and we get $2\times 2$ matrices over $\mathbb{F}_{p}$. In conclusion, modulo evaluation of the entries, the non-injective encoding of the mathematical proof in a sequence of matrices is given by:
$\varphi\leadsto[\varphi]\in M_{2\times 2}(\mathbb{Z}[X_{1},X_{2},\dots])\leadsto[[\varphi]]\in M_{2\times 2}(\mathbb{F}_{p}).$
This leads to the following primitive Zero Knowledge Proof procedure:
  • Consider a mathematical proof $P$, with conclusion $\psi$.
  • Choose values $X_{1},X_{2},\dots\in\mathbb{F}_{p}$.
  • Compute $\alpha_{1}=[[\psi]]$ using the encoding rules.
  • For all axioms $a$ occurring in $P$, compute the corresponding fingerprints $V(a)$ using the encoding rules.
  • For all formulas $\varphi$ occurring in $P$, which are not axioms, compute $V(\varphi)$ by using the homomorphic properties $MP(a,b)$ respectively $Subst_{y}(a,b)$ for appropriated choices of $y,a,b$.
  • The last of these computations produces $\alpha_{2}=[[\psi]]$. Observe that the method to compute $\alpha_{2}$ differs from the method to compute $\alpha_{1}$. While $\alpha_{1}$ was computed by directly encoding $\psi$, $\alpha_{2}$ was computed starting from the axioms and following the homomorphic properties of the proof steps.
  • Check whether $\alpha_{1}=\alpha_{2}$ and accept if this is true, respectively reject, if not.
This Zero Knowledge procedure works grace to the Theorem of Schwartz [4] and Zippel [6]:
Theorem 2.1.
Let $\mathbb{F}$ be a finite field and let $f\in\mathbb{F}[x_{1},\dots,x_{n}]$ be a non-zero polynomial of degree $d\geq 0$. If $r_{1},r_{2},\dots,r_{n}$ are selected randomly and if the choices are independent in $\mathbb{F}$, then:
$Pr[f(r_{1},r_{2},\dots,r_{n})=0]\leq\frac{d}{|\mathbb{F}|}.$
The Schwartz and Zippel Theorem says that for a polynomial which is not identic zero, the probability that a random evaluation is zero in a big finite field is reasonably small. Let again $\varphi$ be the conclusion of the mathematical proof. Let $F_{1}\in M_{2\times 2}(\mathbb{Z}[X_{1},X_{2},\dots])$ the polynomial matrix obtained from $\varphi$ by direct encoding. Let $F_{2}\in M_{2\times 2}(\mathbb{Z}[X_{1},X_{2},\dots])$ be the polynomial matrix obtained from the encoding of the axioms and using the proof steps and their homomorphic properties. We want to prove with high confidence that $F_{1}=F_{2}$ as multivariate polynomials. To this sake, we just compute their evaluations $\alpha_{1}$ and $\alpha_{2}$, where $\alpha_{2}$ is computed from evaluations of the axioms and using homomorphic properties. If the field is sufficiently large, the equality $\alpha_{1}=\alpha_{2}$ means that, with high probability, $F_{1}=F_{2}$, so the formula resulted from the proof is indeed identical with the claimed conclusion, which has been directly encoded.
This primitive Zero Knowledge procedure has the disadvantage that its length is equal with the length of the proof. This procedure must be combined with zero-knowledge proof methods for arithmetic circuits, folding methods, etc, in order to produce ZK-certificates of constant length. Also, the whole procedure must be transformed to a non-interactive one, by applying the Fiat-Shamir Heuristic, [3].
3 Matrices of multivariate polynomials
The original observation which led to this subsection was that matrices consisting of different variables:
$A(k)=\begin{pmatrix}x_{4k+1}&x_{4k+2}\\ x_{4k+3}&x_{4k+4}\end{pmatrix}$
are non-commutative to such extent that if two products are equal:
$A(i_{1})\dots A(i_{n})=A(j_{1})\dots A(j_{m})$
then $n=m$ and $i_{1}=j_{1}$, $\dots$, $i_{n}=j_{n}$. This means that such a monomial (product of elementary matrices) contains information about the number of factors, their order, and their identities. All three elements of information are essential to encode a path inside a tree.
However, proceeding with such matrices would be expensive because one has to choose four field elements to evaluate every elementary matrix and to keep four elements for every matrix to be kept as part of a fingerprint. Instead, we present a system of elementary matrices that achieves the same goal, but needs just one field element to be evaluated and only three field elements for every matrix which is part of a fingerprint.
Definition 3.1.
Let $x_{1}$ be a variable, that is, some element which is transcendental over $\mathbb{Q}$. Let:
$A(x_{1})=\begin{pmatrix}x_{1}&1\\ 0&1\end{pmatrix}$
We consider that $A(x_{1})\in M_{2\times 2}(\mathbb{Z}[x_{1},x_{2},\dots])$.
We show now that these elementary matrices fulfill the same property: a monomial contains enough information to uniquely determine the number of elementary matrices, their order and their identities.
Lemma 3.1.
Consider a set of different variables $V=\{x_{1},x_{2},\dots,x_{k}\}$. Suppose that $0\leq i_{1},\dots,i_{n},j_{1},\dots,j_{m}\leq k$. If:
$A(x_{i_{1}})A(x_{i_{2}})\dots A(x_{i_{n}})=A(x_{j_{1}})A(x_{j_{2}})\dots A(x_{j_{m}}),$
then the following equalities take place: $n=m$, $i_{1}=j_{1}$, $\dots$, $i_{n}=j_{n}$.
Demonstraţie.
If in this identity, we set $x_{1}=\dots=x_{k}=1$, as we observe that:
$\begin{pmatrix}1&n-1\\ 0&1\end{pmatrix}\begin{pmatrix}1&1\\ 0&1\end{pmatrix}=\begin{pmatrix}1&n\\ 0&1\end{pmatrix},$
we get that $n=m$. Let us denote with $S(n)$ the statement: If
$A(x_{i_{1}})A(x_{i_{2}})\dots A(x_{i_{n}})=A(x_{j_{1}})A(x_{j_{2}})\dots A(x_{j_{n}})$
then $i_{1}=j_{1}$, $\dots$, $i_{n}=j_{n}$. We observe that $S(1)$ is evident by identifying the entries. Suppose that we have proved $S(n)$. We look at the hypothesis of $S(n+1)$:
$A(x_{i_{1}})A(x_{i_{2}})\dots A(x_{i_{n+1}})=A(x_{j_{1}})A(x_{j_{2}})\dots A(x_{j_{n+1}}).$
We observe by induction that:
$A(x_{i_{1}})A(x_{i_{2}})\dots A(x_{i_{n}})=\begin{pmatrix}x_{i_{1}}x_{i_{2}}\dots x_{i_{n}}&P(x_{i_{1}},x_{i_{2}},\dots,x_{i_{n-1}})\\ 0&1\end{pmatrix},$
where $P(z_{1},\dots,z_{n-1})\in\mathbb{Z}[z_{1},\dots,z_{n-1}]$ is a fixed polynomial, with the property that no variable $z_{i}$ divides $P$. We write the hypothesis of induction in the form:
$\begin{pmatrix}x_{i_{1}}&1\\ 0&1\end{pmatrix}\begin{pmatrix}x_{i_{2}}x_{i_{3}}\dots x_{i_{n+1}}&P(x_{i_{2}},x_{i_{3}},\dots,x_{i_{n}})\\ 0&1\end{pmatrix}=$
$=\begin{pmatrix}x_{j_{1}}&1\\ 0&1\end{pmatrix}\begin{pmatrix}x_{j_{2}}x_{j_{3}}\dots x_{j_{n+1}}&P(x_{j_{2}},x_{j_{3}},\dots,x_{j_{n}})\\ 0&1\end{pmatrix},$
so:
$\begin{pmatrix}x_{i_{1}}x_{i_{2}}x_{i_{3}}\dots x_{i_{n+1}}&1+x_{i_{1}}P(x_{i_{2}},x_{i_{3}},\dots,x_{i_{n}})\\ 0&1\end{pmatrix}=$
$=\begin{pmatrix}x_{j_{1}}x_{j_{2}}x_{j_{3}}\dots x_{j_{n+1}}&1+x_{j_{1}}P(x_{j_{2}},x_{j_{3}},\dots,x_{j_{n}})\\ 0&1\end{pmatrix}.$
We identify the corresponding entries:
$x_{i_{1}}x_{i_{2}}x_{i_{3}}\dots x_{i_{n+1}}$ $=$ $x_{j_{1}}x_{j_{2}}x_{j_{3}}\dots x_{j_{n+1}},$
$1+x_{i_{1}}P(x_{i_{2}},x_{i_{3}},\dots,x_{i_{n}})$ $=$ $1+x_{j_{1}}P(x_{j_{2}},x_{j_{3}},\dots,x_{j_{n}}).$
We apply the property that no variable $z_{i}$ divides $P$. By variable identification we get $x_{i_{1}}=x_{j_{1}}$. We multiply the hypothesis with $A(x_{i_{1}})^{-1}$ from the left-hand side. We get an instance of $S(n)$ and we apply the induction hypothesis. Of course $A(x_{i_{1}})^{-1}$ does not belong to the matrices over polynomials, but to the matrices over rational functions. It is just important that one can simplify with $A(x_{i})$. ∎
4 Non-commutative edge variables in trees
The goal of this section is to show how to further associate a matrix to a formula represented by a tree using the previous construction.
Below, by edge variable we understand an elementary matrix of the shape:
$X_{i}=\begin{pmatrix}x_{i}&1\\ 0&1\end{pmatrix}.$
To every edge of the tree, and to every vertex, we will associate such a matrix.
In order to represent formulas by trees, both logical and term-building operations are represented as vertices of the tree. For every specific symbol $c$ of arity $d=d(c)$ a number of $d+1$ different edge variables $C,C_{1},\dots,C_{d}$ are associated.
Suppose that a tree $T$ has root $c$ and the sub-trees connected with $c$ are $T_{1},\dots,T_{d}$. Suppose that one already associated matrices
$[T_{1}],\dots,[T_{d}]\in M_{2\times 2}(\mathbb{Z}[X_{1},X_{2},\dots])$
with these sub-trees. Then we associate with $T$ the pair:
$[T]=A(C)+A(C_{1})[T_{1}]+\dots+A(C_{d})[T_{d}].$
Definition 4.1.
If $\varphi$ is a formula or a term, let $[\varphi]$ denote the polynomial matrix associated with its tree.
Theorem 4.1.
A matrix represents at most one formula.
Proof: We show this working out a simplistic example. Consider the following inductive definition:
  1. 1.
    The letters $x$, $y$, $z$ are atomic propositional formulas.
  2. 2.
    If $\varphi$ and $\psi$ are formulas, then:
    $\neg\,\varphi,\,\,\varphi\rightarrow\psi,$
    are formulas.
The alphabet is $A=\{x,y,z,\neg,\rightarrow\}$.
The variables $x,y,z$ are symbols of arity 00 and will always be final nodes. We associate them with the matrices:
$[x]=A(X)=\begin{pmatrix}X&1\\ 0&1\end{pmatrix},\,\,\,\,[y]=A(Y)=\begin{pmatrix}Y&1\\ 0&1\end{pmatrix},\,\,\,\,[z]=A(Z)=\begin{pmatrix}Z&1\\ 0&1\end{pmatrix}.$
[…]
$A(N)=\begin{pmatrix}N&1\\ 0&1\end{pmatrix},\,\,\,\,A(N_{1})=\begin{pmatrix}N_{1}&1\\ 0&1\end{pmatrix}.$
We associate with $\rightarrow$ the matrices:
$A(I)=\begin{pmatrix}I&1\\ 0&1\end{pmatrix},\,\,\,\,A(I_{1})=\begin{pmatrix}I_{1}&1\\ 0&1\end{pmatrix},\,\,\,\,A(I_{2})=\begin{pmatrix}I_{2}&1\\ 0&1\end{pmatrix}.$
The $7$ variables $X,Y,Z,N,N_{1},I,I_{1},I_{2}$ are pairwise different.
The inductive steps are given by:
[$[\neg\,\alpha]=A(N)+A(N_{1})[\alpha],$
$[\alpha\rightarrow\beta]=A(I)+A(I_{1})[\alpha]+A(I_{2})[\beta],$
The statement of the Theorem is proved by induction over the building rules for formulas. What we really prove is the equivalent statement: Every formula is encoded by only one matrix of polynomials. If $\varphi$ is an atomic propositional symbol, then $[\varphi]$ is $[x]$, $[y]$ or $[z]$ and so from $[\varphi]=[\varphi^{\prime}]$ follows immediately $\varphi=\varphi^{\prime}$. Suppose that $\varphi=\neg\,\alpha$. Then:
$[\varphi]=A(N)+A(N_{1})[\alpha].$
We observe that $A(N)$ is the only one monomial of degree one present here. So one can conclude that we are reading a negation. All other monomials start with $A(N_{1})$ because, as shown in Lemma 3.1, all these non-commutative monomials can start only with $A(N_{1})$. Now, by the induction hypothesis, the formula $\alpha$ is uniquely encoded by $[\alpha]$, and it follows that $\varphi$ is uniquely encoded by $[\varphi]$.
Now we consider the case $\varphi=\alpha\rightarrow\beta$. We have seen that:
$[\varphi]=A(I)+A(I_{1})[\alpha]+A(I_{2})[\beta].$
Again $A(I)$ is the only one monomial of degree one and its presence shows that we are reading an implication. All other monomials have the shape $A(I_{1})B$ or $A(I_{2})B$. By the unicity of products of elementary matrices (Lemma 3.1), this monomials can start only with $A(I_{1})$ or with $A(I_{2})$. By common factor, we get the expression $A(I_{1})[\alpha]+A(I_{2})[\beta]$. As by induction hypothesis the formulas $\alpha$ and $\beta$ are uniquely expressed by the polynomial matrices $[\alpha]$, respectively $[\beta]$, it follows that $\varphi$ is uniquely expressed by the matrix of polynomials $[\varphi]$. ∎
5 Homomorphic properties
In this subsection we define the notion of fingerprint of a formula and we show that this notion enjoys homomorphic properties of the operations with formulas used in proofs: modus ponens and substitution. We show how the fingerprint of a formula produced by modus ponens can be computed from the fingerprints of the arguments of the operation modus ponens. Also we show how the fingerprint of a formula obtained by substitution can be computed from the fingerprints of the formula in which the variable has been substituted and the fingerprint of the formula (or term) which has substituted this variable.
Suppose that three different lines of a proof read:
$\varphi\,\,\,\,\,\,\,\,\,\,\,\,\,\,$
$\varphi\rightarrow\psi$
$\psi$
such that the formula $\psi$ is deduced from the formulae $\varphi$ and $\varphi\rightarrow\psi$ by modus ponens. Suppose that we equip the implication symbol $\rightarrow$ with three matrices $A(I)$, $A(I_{1})$ and $A(I_{2})$ such that:
$[\varphi\rightarrow\psi]=A(I)+A(I_{1})[\varphi]+A(I_{2})[\psi].$
Then one can compute the conclusion as follows:
$[\psi]={A(I_{2})}^{-1}\left([\varphi\rightarrow\psi]-A(I)-A(I_{1})[\varphi]\right).$
Substitution also enjoys a homomorphic property. Suppose that one has a formula $\varphi(x)$ and substitutes $x$ with a tree $[\psi]$ corresponding to a formula or a term. We observe that:
$[\varphi(x)]=\sum_{\text{nodes }c}A(X_{i_{1}})\dots A(X_{i_{n}})\cdot A(X_{c}).$
Here for every node $c$, the monomial $A(X_{i_{1}})\dots A(X_{i_{n}})$ consists of the edge-variables on the path from the root to the node $c$. If two such nodes are marked with $x$ and are to be substituted, one has:
$[\varphi(x/\psi)]=[\varphi(x)]-A(X_{i_{1}})\dots A(X_{i_{n}})[x]-A(X_{j_{1}})\dots A(X_{j_{m}})[x]+$
$+A(X_{i_{1}})\dots A(X_{i_{n}})[\psi]+A(X_{j_{1}})\dots A(X_{j_{m}})[\psi].$
In general, let $x\in A$ be a variable, and let $X$ be the polynomial variable associated to this symbol of arity 00. Let $\varphi$ be a formula or a term. We denote by:
$\sum_{c=x}A(X_{i_{1}})\dots A(X_{i_{n}})\cdot A(X_{c}):=[\varphi]_{x}\cdot A(X_{c}).$
It follows that in general for every formula or term $\psi$,
$[\varphi(x/\psi)]=[\varphi]-[\varphi]_{x}\cdot A(X)+[\varphi]_{x}\cdot[\psi].$
Observe that the matrix $[\varphi]_{x}$ has been implicitly defined in the precedent formula.
Definition 5.1.
Let $\varphi$ be a well-formed expression over $A$, i.e. a term or a formula. Suppose that $x_{1},\dots,x_{k}$ are the free variables in $\varphi$. We call the fingerprint of $\varphi$ the tuple:
$([\varphi],[\varphi]_{x_{1}},\dots,[\varphi]_{x_{k}}).$
We denote the fingerprint of $\varphi$ with $F(\varphi)$.
Observe that, as we announced in Section 2, we will deal with two kinds of fingerprints. To a string $\varphi$ which might be a formula or a term, we have defined the fingerprint consisting of a vector of matrices over the polynomial ring $\mathbb{Z}[X_{1},X_{2},\dots]$. Once we fix elements of the finite field for the variables, say $X_{1}=r_{1}$, $\dots$, $X_{k}=r_{k}$, the fingerprint can be evaluated in these values, and becomes a vector of matrices over the finite field. Because of the homomorphic properties presented below, the algebraic relations between polynomial fingerprints are the same as the algebraic relations between evaluated fingerprints.
In the next two theorems we show that the fingerprints of the results of Modus Ponens and substitution can be computed using the fingerprints of the inputs. The proofs are simple computations and we omit them here.
Theorem 5.1.
Suppose that formulas $\varphi$ and $\varphi\rightarrow\psi$ have fingerprints:
$F(\varphi)=([\varphi],[\varphi]_{x_{1}},\dots,[\varphi]_{x_{k}}),$
$F(\varphi\rightarrow\psi)=([\varphi\rightarrow\psi],[\varphi\rightarrow\psi]_{x_{1}},\dots,[\varphi\rightarrow\psi]_{x_{k}}).$
Then the fingerprint of $\psi$ is:
$F(\psi)=([\psi],[\psi]_{x_{1}},\dots,[\psi]_{x_{k}}),$
where:
$[\psi]={A(I_{2})}^{-1}\left([\varphi\rightarrow\psi]-A(I)-A(I_{1})[\varphi]\right),$
$[\psi]_{x_{i}}={A(I_{2})}^{-1}\left([\varphi\rightarrow\psi]_{x_{i}}-A(I_{1})[\varphi]_{x_{i}}\right).$
Theorem 5.2.
Let $\varphi$ and $\psi$ be formulas or terms. Suppose that their fingerprints are:
$F(\varphi)=([\varphi],[\varphi]_{x_{1}},\dots,[\varphi]_{x_{k}}),$
$F(\psi)=([\psi],[\psi]_{x_{1}},\dots,[\psi]_{x_{k}}).$
Let $\varphi(x_{i}/\psi)$ be the result of the substitution of $x_{i}$ with $\psi$ and let $X_{i}$ be a polynomial variable such that $A(X_{i})$ is associated with the $x_{i}$-nodes. Then the expression:
$F(\varphi(x_{i}/\psi))=([\varphi(x_{i}/\psi)],[\varphi(x_{i}/\psi)]_{x_{1}},\dots,[\varphi(x_{i}/\psi)]_{x_{k}})$
where
$[\varphi(x_{i}/\psi)]=[\varphi]-[\varphi]_{x_{i}}\cdot A(X_{i})+[\varphi]_{x_{i}}\cdot[\psi],$
and, if $j\neq i$, then:
$[\varphi(x_{i}/\psi)]_{x_{j}}=[\varphi]_{x_{j}}+[\varphi]_{x_{i}}[\psi]_{x_{j}}$
while if $j=i$, then:
$[\varphi(x_{i}/\psi)]_{x_{i}}=[\varphi]_{x_{i}}[\psi]_{x_{i}}$
Example of a fingerprint: Consider the formula
$\varphi=(x\rightarrow y)\rightarrow(x\rightarrow z).$
According to our definition, one has:
$[\varphi]=[(x\rightarrow y)\rightarrow(x\rightarrow z)]=A(I)+A(I_{1})[x\rightarrow y]+A(I_{2})[x\rightarrow z]=$
$=A(I)+A(I_{1})(A(I)+A(I_{1})A(X)+A(I_{2})A(Y))+$
$+A(I_{2})(A(I)+A(I_{1})A(X)+A(I_{2})A(Z))=$
$=A(I)+A(I_{1})A(I)+A(I_{1})^{2}A(X)+$
$+A(I_{1})A(I_{2})A(Y)+A(I_{2})A(I)+A(I_{2})A(I_{1})A(X)+A(I_{2})^{2}A(Z).$
Also, one has:
$[\varphi]_{x}=A(I_{1})^{2}+A(I_{2})A(I_{1}),$
$[\varphi]_{y}=A(I_{1})A(I_{2}),$
$[\varphi]_{z}=A(I_{2})^{2}.$
Finally, the fingerprint of $\varphi$ is:
$F(\varphi)=([\varphi],[\varphi]_{x},[\varphi]_{y},[\varphi]_{z}).$
[…]
6 Example of a formalized proof
In this section we will present the effort to crystallize a block model proof containing fingerprinting. We consider the following framework of propositional logic. There are three axioms as follows:
$K(\alpha,\beta):\,\,\,\,\alpha\rightarrow(\beta\rightarrow\alpha),$
$S(\alpha,\beta,\gamma):\,\,\,\,(\alpha\rightarrow(\beta\rightarrow\gamma))\rightarrow((\alpha\rightarrow\beta)\rightarrow(\alpha\rightarrow\gamma)),$
$N(\alpha,\beta):\,\,\,\,(\neg\alpha\rightarrow\neg\beta)\rightarrow(\beta\rightarrow\alpha).$
We consider the following theorem:
Theorem 6.1.
$A\rightarrow A.$
The theorem will be proved in three versions. First in classic style. Second, we translate the proof in fingerprinting computations. Third, we represent these computations as blocks.
Classic Proof: Consider the formula $B:=A\rightarrow A$. By making corresponding substitutions, we write down:
$S(A,B,A):\,\,\,\,(A\rightarrow(B\rightarrow A))\rightarrow((A\rightarrow B)\rightarrow(A\rightarrow A)),$
$K(A,B):\,\,\,\,A\rightarrow(B\rightarrow A),$
$K(A,A):\,\,\,\,A\rightarrow(A\rightarrow A),$
At this point we observe that $K(A,A)$ is in fact:
$K(A,A):\,\,\,\,A\rightarrow B,$
$MP(K(A,B),S(A,B,A))=C:\,\,\,\,(A\rightarrow B)\rightarrow(A\rightarrow A),$
$MP(K(A,A),C):\,\,\,\,A\rightarrow A.$
We also observe that the conclusion is the same as $B$. ∎
Fingerprint translation: We consider numeric variables $a,i,i_{1},i_{2}$. In computations, they will be replaced by fixed choices of field elements. The corresponding matrices are:
$A=\begin{pmatrix}a&1\\ 0&1\end{pmatrix},\,\,\,\,I=\begin{pmatrix}i&1\\ 0&1\end{pmatrix},\,\,\,\,I_{1}=\begin{pmatrix}i_{1}&1\\ 0&1\end{pmatrix},\,\,\,\,I_{2}=\begin{pmatrix}i_{2}&1\\ 0&1\end{pmatrix},$
$I_{2}^{-1}=\begin{pmatrix}i_{2}^{-1}&-i_{2}^{-1}\\ 0&1\end{pmatrix},$
Further we define:
$[A]=A,$
$[B]=I+I_{1}[A]+I_{2}[A],$
$[B\rightarrow A]=I+I_{1}[B]+I_{2}[A],$
$[K(A,B)]=I+I_{1}[A]+I_{2}[B\rightarrow A],$
$[K(A,A)]=I+I_{1}[A]+I_{2}[B],$
$[C]=I+I_{1}[K(A,A)]+I_{2}[B],$
$[S(A,B,A)]=I+I_{1}[K(A,B)]+I_{2}[C],$
$I_{2}^{-1}([S(A,B,A)]-I-I_{1}[K(A,B)])=[C],$
$I_{2}^{-1}([C]-I-I_{1}[K(A,A)])=[B].$
Acknowledgment
The author had many discussions on this topic with Grigore Roșu, Xiaohong Chen, Brandon Moore, Miruna Roșca and Traian Șerbănuță. This research was partially supported by a sponsorship from Pi Squared Inc.
Bibliografie
  • [1] Manuel Blum How to Prove a Theorem So No One Else Can Claim It, Proceedings of the International Congress of Mathematicians, Berkeley, California, USA, 1986.
  • [2] Marta Bellés-Muñoz, Miguel Isabel, Jose Luis Muñoz-Tapia, Albert Rubio, Jordi Baylina Circom: A Circuit Description Language for Building Zero-Knowledge Applications, IEEE Transactions on Dependable and Secure Computing, Volume 20, Issue: 6, Nov.-Dec., 2022. DOI: 10.1109/TDSC.2022.3232813
  • [3] Amos Fiat, Adi Shamir How to Prove Yourself: Practical Solutions to Identification and Signature Problems Advances in Cryptology — CRYPTO’ 86. Lecture Notes in Computer Science. Vol. 263. Springer Berlin Heidelberg. pp. 186–194, 1987 DOI:10.1007/3-540-47721-7_12 ISBN 978-3-540-18047-0
  • [4] Jacob T. Schwartz Fast probabilistic algorithms for verification of polynomial identities, Journal of the ACM. 27 (4): 701–717. 1980 DOI: 10.1145/322217.322225
  • [5] Pi Squared Inc. The White Paper, https://pi2.network/papers/pi2-whitepaper, 2025.
  • [6] Richard Zippel Probabilistic algorithms for sparse polynomials, In Ng, Edward W. (ed.) Symbolic and Algebraic Computation, EUROSAM ’79, An International Symposiumon Symbolic and Algebraic Computation, Marseille, France, June 1979, Lecture Notes in Computer Science. Vol. 72. Springer. pp. 216–226. DOI:10.1007/3-540-09519-5_73 ISBN 978-3-540-09519-4.